# Towards Linear Spanners in All Temporal Cliques

### v1

@@ Abstract
Many real-world networks, like transportation networks and social networks, are dynamic in the sense that the edge set may change over time, but these changes are known in advance. This behavior is captured by the temporal graphs model, which has recently become a trending topic in theoretical computer science. A core open problem in the field is to prove the existence of linear-size temporal spanners in temporal cliques, i.e., sparse subgraphs of complete temporal graphs that ensure all-pairs reachability via temporal paths. So far, the best known result is the existence of temporal spanners with $\operatorname{\mathcal{O}}(n\log n)$ many edges. We present significant progress towards proving that linear-size temporal spanners exist in all temporal cliques.
We adapt techniques used in previous works and heavily expand and generalize them to provide a simpler and more intuitive proof of the $\operatorname{\mathcal{O}}(n\log n)$ bound. Moreover, we use our novel approach to show that a large class of temporal cliques, called edge-pivot graphs, admit linear-size temporal spanners. To contrast this, we investigate other classes of temporal cliques that do not belong to the class of edge-pivot graphs. We introduce two such graph classes and we develop novel techniques for establishing the existence of linear temporal spanners in these graph classes as well.
@@ 1 Introduction
Many real-world networks, like transportation networks with scheduled train connections or social networks with repeating meeting schedules, are dynamic, i.e., their edge-set can change over time. To address this, recent research in theoretical computer science has shifted from static graphs to temporal graphs [15, 18]. In these graphs it is assumed that the edge-set can change in every time step but the schedule of the availability of edges is known in advance. Various problems have been studied in this model, including network redesign [10, 12, 13], vertex cover [1, 11, 14], and influence maximization [9]. A general theme is that insights from the static graph setting are only of limited value in temporal graphs, since the temporal availability of the edges makes these problems much harder or even infeasible to solve. For example, even fundamental theorems like Menger’s Theorem do not hold for temporal graphs [17]. Also, one of the most fundamental problems in graphs, that of finding a sparse spanner, is much harder in temporal graphs. A sparse spanner is a small subset of the edge-set that ensures all-pairs connectivity. In static graphs, it is well-known that linear-size spanners exist and can be computed efficiently, e.g., minimum spanning trees [8].
Finding sparse temporal spanners is much more intricate in temporal graphs, since there paths do not necessarily compose. More specifically, Kempe, Kleinberg, and Kumar [15] showed that there is a class of temporal graphs having $\Omega(n\log n)$ many edges, such that no single edge can be removed while preserving temporal connectivity. In the same work, the authors asked, what is the minimum number of edges $c(n)$, such that any temporal graph with $n$ vertices has a temporal spanner using at most $c(n)$ many edges. Almost 15 years later, this question was answered by Axiotis and Fotakis [2] by providing a class of dense but non-complete temporal graphs having $\Theta(n^{2})$ edges, such that any subgraph preserving temporal connectivity necessarily needs to contain $\Omega(n^{2})$ edges as well.
@@
In this paper, we present significant progress on the long-standing open problem by showing that a large class of temporal cliques admits linear-size temporal spanners.
@@ 1.1 Our Contribution
In the first part of the paper, we adapt techniques used in previous works and heavily expand and generalize them. In particular, in [6] a temporal spanner in cliques is computed by a reduction from a clique to a bi-clique. We deepen this connection by showing that the existence of a linear spanner in cliques and bi-cliques is equivalent. Furthermore, in the reduction of Casteigts, Peters, and Schoeters [6], the authors exploit the property that the set containing the earliest edge of each vertex is a perfect matching of the bi-clique and that this also holds for the set of latest edges. We show that this property can also be assumed when searching for bi-spanners in bi-cliques directly. This yields a simpler and more intuitive proof, that temporal cliques admit temporal spanners of size $O(n\log n)$.
Moreover, we reconsider the notion of a pivot-vertex [6, 7]. This is a vertex $u$ that can be reached by every vertex in the graph until some time step $t$, and then $u$ can reach every other vertex in the graph. We transfer this notion to bi-cliques, extend it to pivot-edges, and we strengthen it significantly to find small spanners. If a bi-clique contains a pivot-edge, we can iteratively reduce the size of the graph, while maintaining a linear-size temporal spanner, until the instance is solved, or the instance no longer has a pivot-edge. This shows that a large class of graphs admits linear-size temporal spanners.
In the second part of the paper, we investigate temporal cliques that are not edge-pivot graphs. We found two such graph classes, called shifted matching graphs and product graphs, and for both we provide novel techniques that allow us to show that linear-size spanners always exist. Thus we extend the class of temporal cliques that admit linear-size temporal spanners even further. It is open if our classes cover every temporal clique. However, finding other graphs that do not belong to our graph classes seems to be challenging.
@@ 1.2 Detailed Discussion of Related Work
The study of temporal spanners was initiated by Kempe, Kleinberg and Kumar [15]. In their seminal paper, they introduced the temporal graphs model and showed that spanners on temporal hypercubes have $\Omega(n\log n)$ edges. Axiotis and Fotakis [2] improved on this by providing a class of graphs with $\Theta(n^{2})$ edges, such that no edge can be removed while maintaining temporal connectivity. Later, Casteigts, Peters, and Schoeters [6] studied temporal cliques and showed that temporal spanners of size $O(n\log n)$ always exist. Their "fireworks" algorithm combines different techniques, such as dismountability and delegation, to reduce the clique to a bipartite graph, and then compute a spanner on that graph, that can then be translated to a spanner of the original graph.
Casteigts, Raskin, Renken and Zamaraev [7] took a different approach. They considered random temporal graphs, where the host graph is a clique, and each edge is added to the temporal graph with probability $p$. They showed a sharp threshold for the generated temporal graph being temporally connected, and, even more surprisingly, they showed that the generated graph will contain an almost optimal spanner of size $2n+o(n)$. On a different note, Bilò, D’Angelo, Gualà, Leucci, and Rossi [4] studied whether temporal cliques admit small temporal spanners with low stretch. Their main result shows that they can always compute a temporal spanner with stretch $O(\log n)$ and $O(n\log^{2}n)$ edges.
@@ 2 Preliminaries
Let $(V,E)$ be an undirected graph. This together with a labeling function $\operatorname{\lambda}\colon E\to\operatorname{\mathbb{N}}$ forms the temporal graph $G=(V,E,\operatorname{\lambda})$. Semantically, an edge $e\in E$ is present at time $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({e}}\right)$. For a vertex-set $U\subseteq V$, write $G[U]$ for the subgraph of $G$ induced by $U$. For $v\in V$, let $N(v)$ denote the set of neighbors of $v$, that is, all vertices $u$ with $\mathopen{}\mathclose{{}\left\{{v},{u}}\right\}\in E$. For $v\in V$ and $u,u^{\prime}\in N(v)$, we define the order ${u}\preceq_{v}{u^{\prime}}$, if the edge $\mathopen{}\mathclose{{}\left\{{u},{v}}\right\}$ is not later than the edge $\mathopen{}\mathclose{{}\left\{{u^{\prime}},{v}}\right\}$, i.e., if $\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u}},{{v}}}\right\}}}\right)\leq\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{u^{\prime}}},{{v}}}\right\}}}\right)$. We define $\prec_{v}$, $\succeq_{v}$, and $\succ_{v}$ analogously.
@@
The edges with the smallest and largest label take a special role. For $v\in V$, let $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)\coloneqq\operatorname*{arg\,min}_{u\in N(v)}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{u}}}\right\}}}\right)$ be $v$’s earliest neighbor and let $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({v}}\right)\coloneqq\operatorname*{arg\,max}_{u\in N(v)}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{u}}}\right\}}}\right)$ be $v$’s latest neighbor. We do not consider graphs with isolated vertices and we justify in Section 3 that we only use injective labelings, so this is well-defined. Denote the set of all earliest edges as $\operatorname{\pi^{\text{--}}}\coloneqq\{\mathopen{}\mathclose{{}\left\{{v},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)}}\right\}\mid v\in V\}$ and the set of latest edges as $\operatorname{\pi^{+}}\coloneqq\{\mathopen{}\mathclose{{}\left\{{v},{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({v}}\right)}}\right\}\mid v\in V\}$. If both sets $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ form a perfect matching for $V$, then $G$ is extremally matched. Also, for $v,w\in V$, and $e=\mathopen{}\mathclose{{}\left\{{v},{w}}\right\}\in E$, we write $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)\in[{n}]$ for the index of $e$, when ordering all edges incident to $v$ ascendingly by their label (e.g., if $w=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)$, then $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)=1$).
@@
We focus only on temporal graphs where the underlying graph is a clique or a complete bipartite graph. In these cases, the set of edges is self-evident and for cliques we simply omit it and write $C=(V,\operatorname{\lambda})$. We, similarly, denote complete bipartite temporal graphs by a triple $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $A$ and $B$ are the two partitions and $\operatorname{\lambda}\colon{A}\otimes{B}\to\operatorname{\mathbb{N}}$ is the corresponding labeling function. When necessary, we refer to the set of edges as $E(C)\coloneqq{V}\otimes{V}$ or $E(D)\coloneqq{A}\otimes{B}$. We call $C$ a temporal clique and $D$ a temporal bi-clique. Finally, we define a directed analogue of spanners for temporal bi-cliques. We call a set $S\subseteq{A}\otimes{B}$ a bi-spanner for $D$, if every $a\in A$ can reach every $b\in B$ through a temporal path that is contained in $S$.
@@
(b) A bi-clique where $a_{1}$ cannot reach $a_{2}$ after time $2$ or can be reached by $a_{2}$ until time $4$, so $\operatorname{In}_{v}^{3}\cup\operatorname{Out}_{v}^{3}\neq V$.
@@
Note that for the case of bi-cliques considering both endpoints of $e$ is crucial to ensure that every vertex has a direct edge to an endpoint. Otherwise, we simple examples exist having no vertex $v$ with time $t$ with $\operatorname{In}_{v}^{t}\cup\operatorname{Out}_{v}^{t}=A\sqcup B$. See Figure 0(b).
@@
Finally, for $n\in\operatorname{\mathbb{N}}$, let $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that every temporal clique with $n$ vertices admits a spanner with at most $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges. Analogously, let $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that any bi-clique in which both sides have $n$ vertices admits a bi-spanner with at most $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges. We study the open problem whether $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ holds.
@@ 3 Cliques and Bi-Cliques
We start with a technical lemma that allows us to focus on injective labeling functions for the rest of the paper. Note that the statement holds for all temporal graphs.
@@
Let $G=(V,E,\lambda)$ be a temporal graph. Then there is an injective labeling function $\lambda^{\prime}$ s.t. every spanner $S\subseteq E$ for $(V,E,\lambda^{\prime})$ is also a a spanner for the original graph $G$. In case the original graph $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ is a bi-clique, the same holds for a bi-spanner $S$.
@@
We now show that $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. To this end, we take a bi-clique and add edges within the partitions that yield no benefit in the construction of a spanner.
@@
Let $S$ be a spanner of $C$. To show $S\cap({A}\otimes{B})$ is a bi-spanner in $D$, let $a\in A$ and $b\in B$. As $S$ is a spanner, there is a temporal $ab$-path $P\subseteq S$. If $P$ contains any edges within $A$, no edge to $B$ can be used afterward, due to the labeling. Similarly, once the path arrives at $B$, it is already too late to use any edge within $B$. Thus, $P\subseteq{A}\otimes{B}$ and $S\cap({A}\otimes{B})$ is a bi-spanner in $D$.
@@
Let $D=(A,B,\operatorname{\lambda})$ be a temporal bi-clique with $|A|=|B|=n$, such that any minimum bi-spanner $S^{*}$ in $D$ obeys $\mathopen{}\mathclose{{}\left|S^{*}}\right|=\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. By Lemma 3.8, there is a temporal clique $C=(A\sqcup B,\operatorname{\lambda}^{\prime})$ such that any spanner $S$ in $C$ satisfies $\mathopen{}\mathclose{{}\left|S}\right|\geq\mathopen{}\mathclose{{}\left|S^{*}}\right|$. Since the vertex set of $C$ has size $2n$, there is a spanner of size $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)$ in $C$, we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
@@
We only prove the statement for bi-cliques and $\operatorname{\mathcal{D}}$, the proof for cliques and $\operatorname{\mathcal{C}}$ is analogous. Let $n\in\operatorname{\mathbb{N}}$ and let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique of size $n$. To construct a bi-clique of size $n+1$, we clone one vertex on each side. To be exact, take $a\in A$ and $b\in B$ and let
@@
[…]

### v2

# How to Reduce Temporal Cliques to Find Sparse Spanners

@@ Abstract
Many real-world networks, such as transportation or trade networks, are dynamic in the sense that the edge-set may change over time, but these changes are known in advance. This behavior is captured by the temporal graphs model, which has recently become a trending topic in theoretical computer science. A core open problem in the field is to prove the existence of linear-size temporal spanners in temporal cliques, i.e., sparse subgraphs of complete temporal graphs that ensure all-pairs reachability via temporal paths. So far, the best known result is the existence of temporal spanners with $\operatorname{\mathcal{O}}(n\log n)$ many edges. We present significant progress towards proving whether linear-size temporal spanners exist in all temporal cliques.
We adapt techniques used in previous works and heavily expand and generalize them. This allows us to show that the existence of a linear spanner in cliques and bi-cliques is equivalent and using this, we provide a simpler and more intuitive proof of the $\operatorname{\mathcal{O}}(n\log n)$ bound by giving an efficient algorithm for finding linearithmic spanners. Moreover, we use our novel and efficiently computable approach to show that a large class of temporal cliques, called edge-pivotable graphs, admit linear-size temporal spanners. To contrast this, we investigate other classes of temporal cliques that do not belong to the class of edge-pivotable graphs. We introduce two such graph classes and we develop novel algorithmic techniques for establishing the existence of linear temporal spanners in these graph classes as well.
@@ 1 Introduction
Many real-world networks, like transportation networks with scheduled train connections or social networks with repeating meeting schedules, are dynamic, i.e., their edge-set can change over time. To address this, temporal graphs have gained significant attention in recent research in theoretical computer science  [15, 18]. In these graphs it is assumed that the edge-set can change in every time step but the schedule of the availability of edges is known in advance. Various problems have been studied in this model, including network redesign [10, 12, 13], vertex cover [1, 11, 14], and influence maximization [9]. A general theme is that insights from the static graph setting are only of limited value in temporal graphs, since the temporal availability of the edges makes these problems much harder or even infeasible to solve. For example, even fundamental statements like Menger’s Theorem do not hold for temporal graphs [17]. Also, one of the most essential problems in graphs, that of finding a sparse spanner, is much harder in temporal graphs. A sparse spanner is a small subset of the edge-set that ensures all-pairs connectivity. In static graphs, it is well-known that linear-size spanners exist and can be computed efficiently, e.g., minimum spanning trees [8].
Finding sparse temporal spanners is much more intricate in temporal graphs, since there paths do not necessarily compose (i.e., concatenating two paths where the endpoint of the first is the starting point of the second does not necessarily yield a temporal path). More specifically, Kempe, Kleinberg, and Kumar [15] showed that there is a class of temporal graphs having $\Omega(n\log n)$ many edges, such that no single edge can be removed while preserving temporal connectivity. In the same work, the authors asked, what is the minimum number of edges $c(n)$, such that any temporal graph with $n$ vertices has a temporal spanner using at most $c(n)$ many edges. Almost 15 years later, this question was answered by Axiotis and Fotakis [2] by providing a class of dense but non-complete temporal graphs having $\Theta(n^{2})$ edges, such that any subgraph preserving temporal connectivity necessarily needs to contain $\Omega(n^{2})$ edges as well.
@@
In this paper, we present significant progress on the long-standing open problem by showing that a large class of temporal cliques does admit linear-size temporal spanners.
@@ 1.1 Our Contribution
In the first part of the paper, we adapt techniques used in previous works and heavily expand and generalize them. In particular, in [6] a temporal spanner in cliques is computed by a reduction from a clique to a bi-clique. We deepen this connection by showing that the existence of a linear spanner in cliques and bi-cliques actually is equivalent. Furthermore, in the reduction of Casteigts, Peters, and Schoeters [6], the authors exploit the property that the set containing the earliest edge of each vertex is a perfect matching of the bi-clique and that this also holds for the set of latest edges. We show that this property can also be assumed when searching for bi-spanners in bi-cliques directly. This yields a simple and more intuitive algorithm for computing linearithmic spanners, thus proving the existence of size $O(n\log n)$ temporal spanners in temporal cliques.
Moreover, we reconsider the notion of a pivot-vertex [6, 7]. This is a vertex $u$ that can be reached by every vertex in the graph until some time step $t$, and then $u$ can reach every other vertex in the graph by a temporal path starting at or after time $t$. We transfer this notion to bi-cliques, extend it to $c$-pivot-edges, and we strengthen it significantly to find small spanners. If a bi-clique contains a $c$-pivot-edge, we can iteratively reduce the size of the graph, while maintaining a linear-size temporal spanner, until the instance is solved, or the instance no longer has a $c$-pivot-edge. This reduction rule is widely applicable and yields linear-size temporal spanners in many cases.
In the second part of the paper, we investigate temporal cliques that are not edge-pivotable graphs. We identify two such graph classes, called shifted matching graphs and product graphs, and for both we provide novel techniques that allow us to efficiently compute linear-size spanners. Thus, we extend the class of temporal cliques that admit linear-size temporal spanners even further. It is open if our classes cover every temporal clique. However, finding other graphs that do not belong to our graph classes seems to be challenging.
@@ 1.2 Detailed Discussion of Related Work
The study of temporal spanners was initiated by Kempe, Kleinberg and Kumar [15]. In their seminal paper, they introduce the temporal graphs model and show that spanners on temporal hypercubes have $\Omega(n\log n)$ edges. Axiotis and Fotakis [2] improve on this by providing a class of graphs with $\Theta(n^{2})$ edges, such that no edge can be removed while maintaining temporal connectivity. Later, Casteigts, Peters, and Schoeters [6] study temporal cliques and show that temporal spanners of size $O(n\log n)$ always exist. Their "fireworks" algorithm combines different techniques, such as dismountability and delegation, to reduce the clique to a bipartite graph, and then compute a spanner on that graph, that can then be translated to a spanner of the original graph.
Casteigts, Raskin, Renken and Zamaraev [7] take a different approach. They consider random temporal graphs, where the host graph is a clique, and each edge is added to the temporal graph with probability $p$. They show a sharp threshold for the generated temporal graph being temporally connected. Even more surprisingly, they show that as soon as this threshold is reached, the generated graph will contain an almost optimal spanner of size $2n+o(n)$. On a different note, Bilò, D’Angelo, Gualà, Leucci, and Rossi [4] investigate whether temporal cliques admit small temporal spanners with low stretch. Their main result shows that they can always compute a temporal spanner with stretch $O(\log n)$ and $O(n\log^{2}n)$ edges.
@@ 2 Preliminaries
[…]
@@
The edges with the smallest and largest label take a special role. For $v\in V$, let $\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)\coloneqq\operatorname*{arg\,min}_{u\in N(v)}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{u}}}\right\}}}\right)$ be $v$’s earliest neighbor and let $\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({v}}\right)\coloneqq\operatorname*{arg\,max}_{u\in N(v)}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{v}},{{u}}}\right\}}}\right)$ be $v$’s latest neighbor. We do not consider graphs with isolated vertices and we justify in Section 3 that we only use injective labelings, so this is well-defined. Denote the set of all earliest edges as $\operatorname{\pi^{\text{--}}}\coloneqq\{\mathopen{}\mathclose{{}\left\{{v},{\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)}}\right\}\mid v\in V\}$ and the set of latest edges as $\operatorname{\pi^{+}}\coloneqq\{\mathopen{}\mathclose{{}\left\{{v},{\operatorname{\pi^{+}}\mathopen{}\mathclose{{}\left({v}}\right)}}\right\}\mid v\in V\}$. If both sets $\operatorname{\pi^{\text{--}}}$ and $\operatorname{\pi^{+}}$ form a perfect matching for $V$, then $G$ is extremally matched. Also, for $v,w\in V$, and $e=\mathopen{}\mathclose{{}\left\{{v},{w}}\right\}\in E$, we write $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)$ or simply $\sigma_{v}\mathopen{}\mathclose{{}\left({w}}\right)$ for the index of $e$ in $[{n}]$, when ordering all edges incident to $v$ ascendingly by their label (e.g., if $w=\operatorname{\pi^{\text{--}}}\mathopen{}\mathclose{{}\left({v}}\right)$, then $\sigma_{v}\mathopen{}\mathclose{{}\left({e}}\right)=1$).
@@
We focus only on temporal graphs where the underlying graph is a clique or a complete bipartite graph. In these cases, the set of edges is self-evident and for cliques we simply omit it and write $C=(V,\operatorname{\lambda})$. We, similarly, denote complete bipartite temporal graphs by a triple $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$, where $A$ and $B$ are the two parts and $\operatorname{\lambda}\colon{A}\otimes{B}\to\operatorname{\mathbb{N}}$ is the corresponding labeling function. When necessary, we refer to the set of edges as $E(C)\coloneqq{V}\otimes{V}$ or $E(D)\coloneqq{A}\otimes{B}$. We call $C$ a temporal clique and $D$ a temporal bi-clique. Finally, we define an asymmetric analogue of spanners for temporal bi-cliques. We call a set $S\subseteq{A}\otimes{B}$ a bi-spanner for $D$, if every $a\in A$ can reach every $b\in B$ through a temporal path that is contained in $S$. Note that here, only one side needs to reach the other.
@@
(b) A bi-clique where $a_{1}$ cannot reach $a_{2}$ using a path starting after time $2$ and $a_{1}$ can not be reached by $a_{2}$ using a path arriving before $5$, so $\operatorname{In}_{a_{1}}^{3}\cup\operatorname{Out}_{a_{1}}^{3}\neq V$.
@@
Note that for the case of bi-cliques considering both endpoints of $e$ is crucial to ensure that every vertex has a direct edge to an endpoint. Otherwise, there exist simple instances with a vertex $v$ and a time $t$ with $\operatorname{In}_{v}^{t}\cup\operatorname{Out}_{v}^{t}\neq A\sqcup B$. See Figure 1(b).
@@
Finally, for $n\in\operatorname{\mathbb{N}}$, let $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that every temporal clique with $n$ vertices admits a spanner with at most $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges. Analogously, let $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ be the minimum number such that any bi-clique in which both sides have $n$ vertices admits a bi-spanner with at most $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$ edges. We study the open problem whether $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ and whether $\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)\in\operatorname{\mathcal{O}}(n)$ holds.
@@ 3 Cliques and Bi-Cliques
We start with a technical lemma that allows us to focus on globally injective labeling functions for the rest of the paper. This slightly extends a result by [6] which proved the generality of locally injective labelings. Note that the statement holds for all temporal graphs.
@@
Let $G=(V,E,\lambda)$ be a temporal graph. Then there is an injective labeling function $\lambda^{\prime}$ s.t. every spanner $S\subseteq E$ for $(V,E,\lambda^{\prime})$ is also a spanner for the original graph $G$. In case the original graph $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ is a bi-clique, the same holds for a bi-spanner $S$.
@@
We now show that $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. To this end, we take a bi-clique and add edges within the parts that yield no benefit in the construction of a spanner.
@@
Let $S$ be a spanner of $C$. To show $S\cap({A}\otimes{B})$ is a bi-spanner in $D$, let $a\in A$ and $b\in B$. As $S$ is a spanner, there is a temporal $ab$-path $P\subseteq S$. If $P$ contains any edges within $A$, no edge to $B$ can be used afterwards. Similarly, once the path arrives at $B$, it is already too late to use any edge within $B$. Thus, $P\subseteq{A}\otimes{B}$ and $S\cap({A}\otimes{B})$ is a bi-spanner in $D$.
@@
Let $D=(A,B,\operatorname{\lambda})$ be a temporal bi-clique with $|A|=|B|=n$, such that any minimum bi-spanner $S^{*}$ in $D$ obeys $\mathopen{}\mathclose{{}\left|S^{*}}\right|=\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$. By Lemma 3.8, there is a temporal clique $C=(A\sqcup B,\operatorname{\lambda}^{\prime})$ such that any spanner $S$ in $C$ satisfies $\mathopen{}\mathclose{{}\left|S}\right|\geq\mathopen{}\mathclose{{}\left|S^{*}}\right|$. Since the vertex set of $C$ has size $2n$, there is a spanner of size at most $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)$ in $C$, and thus we have $\operatorname{\mathcal{C}}\mathopen{}\mathclose{{}\left({2n}}\right)\geq\operatorname{\mathcal{D}}\mathopen{}\mathclose{{}\left({n}}\right)$.
@@
We only prove the statement for bi-cliques and $\operatorname{\mathcal{D}}$, the proof for cliques and $\operatorname{\mathcal{C}}$ is analogous. Let $n\in\operatorname{\mathbb{N}}$ and let $D=\mathopen{}\mathclose{{}\left({A},{B},{\operatorname{\lambda}}}\right)$ be a bi-clique of size $n$. We construct a bi-clique of size $n+1$ with at least the same minimum spanner size as $D$ by cloning one vertex on each side. To be exact, take $a\in A$ and $b\in B$ and let
@@
$\operatorname{\lambda}^{\prime}$ $\coloneqq x,y\mapsto\begin{cases}\operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{a}},{{b}}}\right\}}}\right),\quad&\text{if }x=a^{\prime},y=b^{\prime},\\ \operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{a}},{{y}}}\right\}}}\right),\quad&\text{else if }x=a^{\prime},\\ \operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{x}},{{b}}}\right\}}}\right),\quad&\text{else if }y=b^{\prime},\\ \operatorname{\lambda}\mathopen{}\mathclose{{}\left({\mathopen{}\mathclose{{}\left\{{{x}},{{y}}}\right\}}}\right),\quad&\text{otherwise},\end{cases}$